# Bayesian Sheaf Neural Networks

### v1

@@ 3.2 Variational Sheaf Learner
Case 3: Invertible Diagonal. In the case where the sheaf restriction maps $\mathcal{F}_{u\trianglelefteq e}:\mathcal{F}(u)\to\mathcal{F}(e)$ are assumed to be invertible diagonal matrices with respect to the standard basis, $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ are $d$--dimensional vectors and we take $q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)=Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ to be a normal distribution with diagonal covariance, so that $\mathcal{F}_{u\trianglelefteq e}=\operatorname{diag}(z)$ for $z\sim\mathcal{N}_{d}(\mu_{u\trianglelefteq e},\operatorname{diag}(\sigma_{u\trianglelefteq e}))$. The reparameterization $\mathcal{F}=g_{\phi}(X,\epsilon)$ in this case is given by setting $\mathcal{F}_{u\trianglelefteq e}=\operatorname{diag}(\mu_{u\trianglelefteq e}+\sigma_{u\trianglelefteq e}\epsilon)$ for $\epsilon\sim\mathcal{N}(\mathbf{0},I_{d})$. Lastly, we use a standard normal prior, where $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ and $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})=\mathcal{N}_{d}(\mathbf{0},I_{d})$.
3.3 Bayesian Sheaf Neural Network Overview
@@
We give an overview of the Bayesian sheaf neural network architecture, which is also summarized in 2. The variational sheaf learning mechanism maps the input node features to the sheaf distributional parameters as described in Section 3.2, though we briefly describe here as well. A linear layer first resizes the input node features to dimension $k$. Specifically, we use $k=df$ where $d$ and $f$ are the stalk dimensions and number of feature channels respectively. Then, for each incident node-edge pair $u\trianglelefteq e$, the features of the pair of nodes corresponding to the edge $e$ are concatenated. Hence a node feature matrix of size $n\times k$ is transformed by this concatenation operation into a matrix of size $2e\times 2k$, where $n$ denotes the number of nodes, and $e$ denotes the number of edges in the graph. Lastly, an MLP maps these features indexed by the incident node-edge pairs $u\trianglelefteq e$ to distributional parameters $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ as described by Eq. (11), which define the distribution over the cellular sheaves. From this distribution, a collection of independent and identically distributed sheaves $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{L}$ are sampled: one for each layer of the network.
Separately, a multi-layer perceptron maps the input node features $X\in\mathbb{R}^{n\times m}$ to features of dimension $X^{\prime}\in\mathbb{R}^{n\times df}$ where $d$ is the stalk dimension and $f$ is the number of hidden channels. The features $X^{\prime}$ are reshaped to size $nd\times f$. Sheaf layers described by Eq. (5) update the transformed features $X^{\prime}$ using the sampled sheaves $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{L}$. Finally, a linear layer maps the output of the final sheaf layer to a vector of class probabilities.
During training, a one-sample Monte Carlo estimate of Eq. (9) or Eq. (10) is computed, and the loss is back-propagated to update the network weights. During testing, ensembling is performed to obtain point estimates for the unobserved node labels $y^{*}$.
4 Reparameterizable Distributions on $SO(n)$
In order to perform variational inference for our BSNN when cellular sheaves with special orthogonal restriction maps are used, we require a family of distributions on $SO(n)$ that are both reparameterizable and have tractable density functions. We define a novel family of such distributions via the Cayley transform [13].
4.1 Cayley Distributions on $SO(n)$
We denote the set of all $n\times n$ skew-symmetric matrices by $\mathfrak{so}(n)$. The Cayley transform is a function $C:\mathfrak{so}(n)\to SO(n)$ defined by
$C(A)=(I-A)^{-1}(I+A)=2(I-A)^{-1}-I.$ (12)
If $A$ is an $n\times n$ skew-symmetric matrix, then $x^{\mathsf{T}}Ax=0$ for all $x\in\mathbb{R}^{n}$ due to the fact that $x^{\mathsf{T}}Ax=(x^{\mathsf{T}}Ax)^{\mathsf{T}}=-x^{\mathsf{T}}Ax$. Hence if $(I-A)x=0$, then $x^{\mathsf{T}}(I-A)x=x^{\mathsf{T}}x=0$ implying that $x=0$. Thus $I-A$ is invertible, showing that the Cayley transform is well-defined. If $P\in SO(n)$ does not have $-1$ as an eigenvalue, $C^{-1}(P)$ exists and is given by
$C^{-1}(P)=(P-I)(I+P)^{-1}=I-2(I+P)^{-1}.$ (13)
Remark 4.1.
For a vector $\phi\in\mathbb{R}^{n(n-1)/2}$, let $X_{\phi}$ denote the skew symmetric matrix containing the entries of $\phi$ as its lower triangular entries. Explicitly, if $\phi=(\phi_{1},\phi_{2},\dots,\phi_{n(n-1)/2})$ and if $x_{ij}$ denotes the $i,j$--th entry of $X_{\phi}$, then for $i>j$, set $x_{ij}=-x_{ji}=\phi_{ij-j(j+1)/2}$ and set $x_{ii}=0$. We may vectorize the Cayley transform and regard it as a map $\widetilde{C}:\mathbb{R}^{n(n-1)/2}\to\mathbb{R}^{n^{2}}$ by setting $\widetilde{C}(\phi)=\operatorname{vec}C(X_{\phi})$. The Jacobian (determinant) of $\widetilde{C}$ is $J\widetilde{C}(\phi)=\det(D\widetilde{C}(\phi)^{\mathsf{T}}D\widetilde{C}(\phi))^{1/2}$, where $D\widetilde{C}(\phi)$ denotes the total derivative of $\widetilde{C}$ at the point $\phi$. The Jacobian $J\widetilde{C}(\phi)$ can be expressed [41] in terms of $X_{\phi}$ by
$J\widetilde{C}(\phi)=\frac{2^{3n(n-1)/4}}{\det(I+X_{\phi})^{n-1}}.$ (14)
Definition 4.2.
Let $C:\mathfrak{so}(n)\to SO(n)$ denote the Cayley transform and let $X$ be a uniformly distributed random variable on $SO(n)$. Given parameters $M\in SO(n)$ and $10\leq\kappa<$ 1, we define the Cayley distribution $\mathcal{C}_{n}(M,\kappa)$ on $SO(n)$ to be the distribution of $Y=C(\frac{1-\kappa}{1+\kappa}C^{-1}(X))M$.
Remark 4.3.
Note that by the definition of the Cayley distribution, a random variable $Y\sim\mathcal{C}_{n}(M,\kappa)$ can be expressed as a deterministic function of $M$, $\kappa$, and $X\sim U_{SO(n)}$. Hence these distributions are reparameterizable. For details on sampling from the uniform distribution on $SO(n)$, see for example [1].
Theorem 4.4.
The Cayley distribution $\mathcal{C}_{n}(M,\kappa)$ has density
with respect to the normalized Haar measure on $SO(n)$.
For a proof see Appendix B.
Remark 4.5.
The family of Cayley distributions defined in [41] have density proportional to $\det(PM^{\mathsf{T}}+I)^{\kappa}$ of Eq. (15). Note that the concentration parameter $\kappa$ appears in the exponent of the determinant expression, whereas in our proposed family of Cayley distributions, $\kappa$ instead scales the identity matrix term within the determinant.
4.2 Special Cases of the Cayley Distribution
We now examine the distributions $\mathcal{C}_{n}(M,\kappa)$ for small values of $n$.
Case $\mathbf{n=3}$. Note that $SO(3)$ is homeomorphic to $3$--dimensional real projective space $RP^{3}$, which is the quotient of $S^{3}$ under the identification of antipodal points. We show that $\mathcal{C}_{3}(M,\kappa)$ is related to an angular central Gaussian distribution [52] on $S^{3}$. Specifically, we show that $\mathcal{C}_{3}(M,\kappa)$ is the pushforward of an angular central Gaussian distribution on $S^{3}$ under the quotient map $\Phi:S^{3}\to SO(3)$. If $x\in\mathbb{R}^{n+1}$ is a unit vector, $\Lambda\in\mathbb{R}^{(n+1)\times(n+1)}$ is a symmetric positive definite matrix, and $\alpha_{n}$ denotes the volume of the $n$--dimensional sphere $S^{n}$, the angular central Gaussian distribution $ACG_{n}(\Lambda)$ on $S^{n}$ has density
$f(x;\Lambda)=\alpha_{n}^{-1}\det(\Lambda)^{-\frac{1}{2}}(x^{\mathsf{T}}\Lambda^{-1}x)^{-2}$
with respect to the uniform measure on $S^{n}$.
Proposition 4.6.
For a proof see Appendix B.
When $n=2$ or $n=3$, Eq. (14) for the Jacobian of the Cayley transform simplifies to $2^{3n(n-1)/4}(1+\|\phi\|^{2})^{1-n}$. One consequence of this fact is that $\mathcal{C}_{n}(I_{n},\kappa)$ is the pushforward of a Cauchy distribution (multivariate in the case $n=3$) under the Cayley transform. This fact makes it possible to compute the KL divergence of $\mathcal{C}_{n}(M,\kappa)$ from the uniform distribution on $SO(n)$ in these cases.
Proposition 4.7.
For a proof see Appendix C.
Thus when learning a distribution of cellular sheaves with special orthogonal restriction maps in our BSNN, for the cases where the stalk dimension $d$ is $2$ or $3$, we are able to use Eq. (9) to estimate the ELBO by plugging in the expressions for the KL divergence from Proposition 4.7. When $d\geq 4$, we estimate the ELBO by simply using Eq. (10).
5 Node Classification Experiments
We compare our Bayesian sheaf neural networks (BSNN) against the sheaf neural networks (NSD) from [6] on the WebKB datasets, a set of real-world web-page datasets. These datasets have low edge-homophily coefficients, that is, there are relatively few edges which connect nodes belonging to the same class [58]. This is a type of graph data where sheaf neural networks are expected to outperform traditional graph neural networks. See Appendix D for further details on the datasets. To study the performance of our Bayesian sheaf models when training data is limited, we use 10 fixed splits of the datasets, each of which contains $32\%/20\%/48\%$ of the nodes for training, validation, and testing respectively. For a given hyperparameter configuration, each model is trained and evaluated on 10 fixed splits of the data and we compute the average validation accuracy. Identical hyperparameter sweeps are performed for each model. For the hyperparameter configuration that yields the best validation accuracy for a model, we train and evaluate the model on the 10 splits for 30 different random seeds. We report the mean test accuracies along with the standard deviation of the test accuracies across the 30 random seeds in Table 1. For each type of restriction map (diagonal, special orthogonal, and general linear), we compare the results of the BSNN model against the NSD model. The higher test accuracy is bolded, and further, a blue entry denotes that a Wilcoxon signed rank test indicates a statistically significant difference for a $p$-value threshold of $p=0.05$. We also include results for a graph convolutional network (GCN) as a baseline comparison.
Additionally, to examine the sensitivity of the Bayesian and deterministic sheaf models to the hyperparameter selections, we report the mean test accuracy across all hyperparameter configurations in Table 2.

### v2

@@ 3.2 Variational Sheaf Learner
Case 2: Invertible Diagonal. In the case where sheaf restriction maps $\mathcal{F}_{u\trianglelefteq e}$ are assumed to be invertible diagonal matrices with respect to the standard basis, $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ are $d$-dimensional vectors and we take $q_{\phi}(\mathcal{F}_{u\trianglelefteq e}|X,y)=Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$ to be a normal distribution with diagonal covariance, so that $\mathcal{F}_{u\trianglelefteq e}=\operatorname{diag}(z)$ for $z\sim\mathcal{N}_{d}(\mu_{u\trianglelefteq e},\operatorname{diag}(\sigma_{u\trianglelefteq e}))$. The reparameterization $\mathcal{F}=g_{\phi}(X,\epsilon)$ in this case is given by setting $\mathcal{F}_{u\trianglelefteq e}=\operatorname{diag}(\mu_{u\trianglelefteq e}+\sigma_{u\trianglelefteq e}\epsilon)$ for $\epsilon\sim\mathcal{N}(\mathbf{0},I_{d})$. Lastly, we use a standard normal prior, where $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ and $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})=\mathcal{N}_{d}(\mathbf{0},I_{d})$.
Case 3: Special Orthogonal. The class of sheaves with special orthogonal restriction maps is of particular interest for sheaf neural networks, including our Bayesian sheaf neural networks, due to both theoretical and practical properties. From a theoretical point of view, certain results concerning the kernel of $\Delta_{\mathcal{F}}$ can be proven for sheaves using rotations as restriction maps [10, Section 3.1]. Regarding the more practical aspects, rotations offer a nice middle-ground between diagonal and general linear maps in terms of balancing the number of parameters with the expressivity of the resulting sheaf. Moreover, when using rotations, the blocks along the diagonal of the sheaf Laplacian $L_{\mathcal{F}}$ are scalar multiples of the identity matrix. This makes computing the normalized sheaf Laplacian $\Delta_{\mathcal{F}}$ easier.
These properties motivate our focus on $SO(d)$ sheaves. To further support their utility, we extend previous theoretical results on their separation power as defined in Section 2.2. In particular, we prove that for any number of classes $C\in\mathbb{N}$, we may find $d$ large enough such that $\mathcal{H}_{so}^{d}$ has linear separation power over $\mathcal{G}_{C}$. In the following proposition, $\lfloor\cdot\rfloor$ denotes the floor function.
Proposition 1.
For a proof see Appendix Appendix A: Proof of Proposition 1.
Remark 6.
An extra consequence of Proposition 1 is that it also provides a new upper bound on the stalk dimension $d$ necessary for $\mathcal{H}^{d}_{gen}$, the collection of sheaves with $d\times d$ general linear restriction maps, to have linear separation power over the collection of connected graph $\mathcal{G}_{C}$ with $C$ classes. This is due to the simple fact that $\mathcal{H}^{d}_{so}\subseteq\mathcal{H}^{d}_{gen}$.
We now turn to the practical challenge of incorporating these $SO(d)$ maps into the variational model. Typical probability distributions on $SO(d)$, such as the matrix Langevin distribution, are not easily reparameterizable. This makes incorporating such distributions into neural networks that are trained via back-propagation difficult. On the other hand, while reparameterizable distributions on $SO(d)$ have been constructed in [27] using the exponential map, the density functions of such distributions can be difficult to compute exactly, which obstructs one from computing, or even efficiently estimating, the KL divergence term in Eq.(18).
In Section 3.3, we introduce Cayley distributions, a family of reparameterizable distributions $\mathcal{C}_{d}(M,\kappa)$ on $SO(d)$ with closed form expressions for their probability density functions. We use $\mathcal{C}_{d}(M,\kappa)$ as the variational distributions for sheaves with special orthogonal restriction maps. Hence when $\mathcal{F}_{u\trianglelefteq e}\in SO(d)$, the distributional parameters are a mean rotation $\mu_{u\trianglelefteq e}\in SO(d)$ and scalar concentration parameter $\sigma_{u\trianglelefteq e}\in[0,1)$, and we have $Q(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})=\mathcal{C}_{d}(\mu_{u\trianglelefteq e},\sigma_{u\trianglelefteq e})$. The reparameterization of the Cayley distributions is described in Remark 7. We use a prior distribution $p_{\theta}(\mathcal{F})=\prod_{u\trianglelefteq e}p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ where each $p_{\theta}(\mathcal{F}_{u\trianglelefteq e})$ is a uniform distribution on $SO(d)$.
3.3 Cayley Distributions on $SO(n)$
To enable reparameterizable inference on $SO(n)$, we define a novel family of distributions using the Cayley transform.
Definition 8.
Let $C:\mathfrak{so}(n)\to SO(n)$ denote the Cayley transform and let $X$ be a uniformly distributed random variable on $SO(n)$. Given parameters $M\in SO(n)$ and $0\leq\kappa<1$, we define the Cayley distribution $\mathcal{C}_{n}(M,\kappa)$ on $SO(n)$ inspired by the constructions in [50] to be the distribution of $Y=C\left(\frac{1-\kappa}{1+\kappa}C^{-1}(X)\right)M$.
Remark 7.
Note that by the definition of the Cayley distribution, a random variable $Y\sim\mathcal{C}_{n}(M,\kappa)$ can be expressed as a deterministic function of $M$, $\kappa$, and $X\sim U_{SO(n)}$. Hence these distributions are theoretically reparameterizable. For details on sampling from the uniform distribution on $SO(n)$, see for example [2].
Theorem 1.
Let $P\in SO(n)$, the Cayley distribution $\mathcal{C}_{n}(M,\kappa)$ has density
with respect to the normalized Haar measure on $SO(n)$.
For a proof see Appendix Appendix B: Section 3.3 Proofs.
Remark 8.
The family of Cayley distributions defined in [50] have density proportional to $\det(PM^{\mathsf{T}}+I)^{\kappa}$ of Eq.(22). Note that the concentration parameter $\kappa$ appears in the exponent of the determinant expression, whereas in our proposed family of Cayley distributions, $\kappa$ instead scales the identity matrix term within the determinant. This structural difference plays a key role in enabling reparameterizable constructions based on the Cayley transform, particularly in low dimensions.
For small values of $n,$ the structure of $SO(n)$ allows for simpler parameterizations and more efficient sampling schemes. We now examine these cases to motivate practical approaches for facilitating differentiable sampling from $\mathcal{C}_{n}(M,\kappa)$.
Case $\mathbf{n=3}$. Note that $SO(3)$ is homeomorphic to 33-dimensional real projective space $RP^{3}$, which is the quotient of $S^{3}$ under the identification of antipodal points. We show that $\mathcal{C}_{3}(M,\kappa)$ is related to an angular central Gaussian distribution [61] on $S^{3}$. Specifically, we show that $\mathcal{C}_{3}(M,\kappa)$ is the pushforward of an angular central Gaussian distribution on $S^{3}$ under the quotient map $\Phi:S^{3}\to SO(3)$. This provides a reparameterizable sampling strategy, distinct from sampling via uniform $SO(3)$.
If $x\in\mathbb{R}^{n+1}$ is a unit vector, $\Lambda\in\mathbb{R}^{(n+1)\times(n+1)}$ is a symmetric positive definite matrix, and $\alpha_{n}$ denotes the volume of the $n$-dimensional sphere $S^{n}$, the angular central Gaussian distribution $ACG_{n}(\Lambda)$ on $S^{n}$ has density
$f(x;\Lambda)=\alpha_{n}^{-1}\det(\Lambda)^{-\frac{1}{2}}(x^{\mathsf{T}}\Lambda^{-1}x)^{-2},$
with respect to the uniform measure on $S^{n}$.
Proposition 2.
For a proof see Appendix Appendix B: Section 3.3 Proofs.
When $n=2$ or $n=3$, Eq.(6) for the Jacobian of the Cayley transform simplifies to $2^{3n(n-1)/4}(1+\|\phi\|^{2})^{1-n}$. One consequence of this fact is that $\mathcal{C}_{n}(I_{n},\kappa)$ is the pushforward of a Cauchy distribution (multivariate in the case $n=3$) under the Cayley transform. This fact makes it possible to compute the KL divergence of $\mathcal{C}_{n}(M,\kappa)$ from the uniform distribution on $SO(n)$ in these cases. Note that while the prior is uniform on $SO(n)$, reparameterizable sampling is achieved via Cauchy or $ACG_{3}$ distributions in latent space.
Proposition 3.
For a proof see Appendix Appendix C: KL Divergence Calculations.
Thus when learning a distribution of cellular sheaves with special orthogonal restriction maps in our BSNN, for the cases where the stalk dimension $d$ is $2$ or $3$, we are able to use Eq.(19) to estimate the ELBO by plugging in the expressions for the KL divergence from Proposition 3. When $d\geq 4$, we estimate the ELBO by simply using Eq.(20).
3.4 Bayesian Sheaf Neural Network Architecture
We present the Bayesian sheaf neural network architecture, which is also summarized in Figure 2 and Algorithm 1. The variational sheaf learning mechanism maps the input node features to the sheaf distributional parameters as described in Section 3.2, For completeness, we briefly review the mechanism here as well.
A linear layer first resizes the input node features to dimension $k$. Specifically, we use $k=df$ where $d$ and $f$ are the stalk dimensions and number of feature channels respectively. For each edge $e=(u,u^{\prime})\in E$, we consider the two incident node-edge pairs $u\trianglelefteq e$ and $u^{\prime}\trianglelefteq e$. For each pair, we construct a feature vector by concatenating the features of the node with those of the node at the other end of the edge. Hence, a node feature matrix of size $n\times k$ is transformed by this concatenation operation into a matrix of size $2e\times 2k$, where $n$ denotes the number of nodes, and $e$ denotes the number of edges in the graph.
Lastly, an MLP maps these features indexed by the incident node-edge pairs $u\trianglelefteq e$ to distributional parameters $\mu_{u\trianglelefteq e}$ and $\sigma_{u\trianglelefteq e}$ as described by Eq.(21), which define the distribution over the cellular sheaves. From this distribution, a collection of independent and identically distributed sheaves $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{L}$ are sampled: one for each layer of the network.
Separately, a multi-layer perceptron maps the input node features $X\in\mathbb{R}^{n\times m}$ to features of dimension $X^{\prime}\in\mathbb{R}^{n\times df}$ where $d$ is the stalk dimension and $f$ is the number of hidden channels. The features $X^{\prime}$ are reshaped to size $nd\times f$.
Sheaf layers described by Eq.(3) update the transformed features $X^{\prime}$ using the sampled sheaves $\mathcal{F}_{1},\mathcal{F}_{2},\dots,\mathcal{F}_{L}$. Finally, a linear layer maps the output of the final sheaf layer to a vector of class probabilities.
During training, a one-sample Monte Carlo estimate of Eq.(19) or Eq.(20) is computed, and the loss is back-propagated to update the network weights. During testing, multiple sheaf samples are drawn from the learned posterior, and ensembling is performed over the corresponding forward passes to obtain point estimates for the unobserved node labels $y^{*}$.
@@ 3.4 Bayesian Sheaf Neural Network Architecture
Algorithm 1 Bayesian Sheaf Neural Network (BSNN)
0:  Node features $X\in\mathbb{R}^{n\times k}$, graph $G=(V,E)$, stalk dim $d$, hidden channels $f$, layers $L$
0:  Predicted class probabilities for nodes
  $k\leftarrow d\cdot f$
  $X\leftarrow\text{Linear}(X)$ {Resize to $n\times k$}
  Initialize $H\leftarrow[\,]$
  for each edge $e=(u,u^{\prime})\in E$ do
   Append $X_{u}\|X_{u^{\prime}}]$ to $H$
   Append $X_{u^{\prime}}\|X_{u}]$ to $H$
  end for
  $(\mu,\sigma)\leftarrow\text{MLP}(H)$
  Sample sheaves $F_{1},\dots,F_{L}$
  $X^{\prime}\leftarrow\text{MLP}(X)$
  Reshape $X^{\prime}$ to $nd\times f$
  for $\ell=1$ to $L$ do
   $X^{\prime}\leftarrow\text{SheafLayer}(X^{\prime},F_{\ell})$
  end for
  $\hat{y}\leftarrow\text{Linear}(X^{\prime})$
  if training then
   Compute ELBO and backpropagate
  else
   Ensemble predictions to estimate $\hat{y}^{*}$
  end if
4 Experiments and Uncertainty Quantification
We compare our proposed model against a comprehensive set of baseline methods, including classical and advanced graph neural networks (GNNs), and modern deep geometric learning architectures.